OpenStudy (anonymous):

solve using the square root method. 5x^2 = 60

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

5x^2 = 60 => x^2=12. so x=+sqrt(12) or -sqrt(12)

OpenStudy (anonymous):

Not entirely sure what you mean by square root method. U have x^2 =12 so x is plus/minus sqrt 12

OpenStudy (anonymous):

okay thank u!